OpenStudy (anonymous):

Find a third-degree polynomial with real coefficients and with zeros –3 and –4 + i.

OpenStudy (anonymous):

one factor is \(x+3\) for sure the hard part is finding a quadrtic with zeros \(-4+i\) and\(-4-i\) there are three ways do to it one is hard one is easy one is real real easy you pick

OpenStudy (misty1212):

lol pick the hard way \(\color\magenta\heartsuit\)

OpenStudy (anonymous):

okay, lol.

OpenStudy (anonymous):

lets do the easy way first ] work backwards starting with \[x=-4+i\] then add 4 to both sides to get \[x+4=i\] square both sides (carefully) to get \[(x+4)^2=-1\] or \[x^2+8x+16=-1\]

OpenStudy (anonymous):

adding 1 gives you the quadratic as \[x^2+8x+17\]

OpenStudy (anonymous):

that is the easy wat

OpenStudy (misty1212):

i guess that was easy...

OpenStudy (anonymous):

the real real easy way requires memorizing something namely, that if \(a+bi\) is a zero, the quadratic is \[x^2-2ax+(a^2+b^2)\] in your case \(a=-4,b=1\) so \[x^2-2(-4)x+(6^2+1^2)\] \[x^2+8x+17\]

OpenStudy (anonymous):

typo there, but you get the idea

OpenStudy (misty1212):

yeah \[x^2-2(-4)x+((-4)^2+1^2)\]

OpenStudy (misty1212):

lot easier than trying \[(x-(-4+i))(x-(-4-i))\]

OpenStudy (anonymous):

final job is to multiply \[(x+3)(x^2+8x+17)\]

OpenStudy (beginnersmind):

\[(x-(-4+i))(x-(-4-i))\] isn't hard either. It's in the form [(x+4)+i] * [(x+4)-i] which equals: (x+4)^2 + 1